Suppose that there are 36 persons in a room, ten of which are females. From among these 36 persons, eight will be choosen to serve on a committee.
a) if there must be four males adn four females on the committee, then how many committes of this type would be possible? Show work.
Answer #1

a) number of such committes =N(select 4 male from 26 male and 4 female from 10)

=(26C4)*(10C4) =14950*210 =3139500